OpenStudy (anonymous):

Solve the equation. 3^(2x-1) . 9^(x+1)= 27^(x)

OpenStudy (hunus):

\[3^{2x-1}* 3^{2(x+1)}=3^{3x}\] \[(2x-1)+(2x+2)=3x\] \[4x+1=3x\] \[x=(-1)\]